OpenStudy (anonymous):

what is y=4/3x-2 3y-4x=-6

OpenStudy (anonymous):

This is an equation of a straight line of the form y = mx + c where the slope (m) = 4/3 and y-intercept (c) = -2. Please checkout the following link for details: http://www.mathsisfun.com/equation_of_line.html